### 08 Jan 2024 v1

@@ 3.2 In-distribution generalization
By in-distribution generalization we refer to GrainGNN inference for $L=L^{0}$, unseen (during training) $\boldsymbol{p}$ values, and unseen $\boldsymbol{\xi}$ values sampled by the same procedure we used for training. We show an example of test case in Fig. 7 with $\boldsymbol{p}=(1.904\ K/\mu m,0.558\ m/s)$. Using nearest neighbor interpolation in $\boldsymbol{p}$ space we determine that GrainGNN should be used with $\Delta z=2.4\mu m$, which determines the number of GrainGNN steps. Fig. 7a shows the grain microstructure at four different heights. At each height $z_{l}$, we show the $G_{l}$ and $I_{l}$ with $500\times 500$ spatial resolution. We compare the reconstructed image with the phase field data at the same height. The misclassified grid points are marked in dark red. The error images indicate that the GrainGNN predictions are accurate representations of the microstructure obtained by our phase field simulations. The last two columns compare the evolution of the 3D microstructure for GrainGNN and phase field predictions. We select two grains to illustrate the grain shape evolution. Fig. 7b-e are different measures of GrainGNN accuracy. Fig. 7b shows the cumulative grain elimination events when the SLI reaches different heights. The blue, red line, dashed red lines depict grain eliminations that happened in the phase field simulation only, GrainGNN inference only, and in both, respectively. We can see that the number of eliminations predicted by GrainGNN is overall accurate but lagged after $z_{l}=24\ \mu m$ compared to the phase field result. At $z_{l}=50\ \mu m$, three grains that should be eliminated still exist on the graph, and only one grain is falsely eliminated by GrainGNN. The precision and recall for grain events are 72/73 and 72/75 respectively. Fig. 7c plots MR for reconstructed images at different heights. MR is 2.4% initially at $z_{l}=2\ \mu m$ because the curvature of the grains is neglected in reconstructed images. MR is 11.3% at $z_{l}=50\ \mu m$. Fig. 7d shows the evolution of the volume-averaged misorientation as the height of the SLI increases; the curve is well captured by GrainGNN. The initial average misorientation angle is $23.8^{\circ}$ and it decreases to $13.8^{\circ}$, which indicates the grains are more aligned with the temperature gradients after the epitaxial growth. Fig. 7e shows the grain size distribution when the SLI reaches $50\mu m$. The final average grain size is $9.9\mu m$. The KS statistic between the phase field and GrainGNN distributions is 0.034.
@@
Fig. 9: Error statistics for in-distribution and out-of-distribution generalization. (a) Black lines are $\boldsymbol{p}$ grid used for training and validation. Black dots are 100 $\boldsymbol{p}$ values for testing. (b) Misclassification rate (MR) averaged across sampled heights $z_{l}$ for each testing $\boldsymbol{p}$. (c) KS statistics (KS) at the end of simulations for testing $\boldsymbol{p}$. (d) MR mean and standard deviation for 100 testing $\boldsymbol{p}$ with domain size $(40\mu m,40\mu m,50\mu m)$. (e) MR for domain size $(120\mu m,120\mu m,50\mu m)$. (f) MR for domain size $(80\mu m,80\mu m,100\mu m)$. (g) MR and KS for different initial mean grain sizes $d_{0}$. (h) MR and KS for different initial grain orientation distributions. $\theta_{0}$ is the most frequent misorientation angle between the grain orientation and the $z$-axis.
@@
We further test GrainGNN’s MR accuracy on 100 randomly selected $\boldsymbol{p}$ with a single random $\boldsymbol{\xi}$ for each value of $\boldsymbol{p}$. (see Fig. 9a ). Fig. 9b show the MR averaged over $z_{l}$ for each testing case. MR increases with increasing $R_{z}$ and decreasing $G_{z}$; the highest MR is 17% for $G_{z}=2.450K/\mu m,R_{z}=1.753m/s$ and $G_{z}=3.764K/\mu m,R_{z}=1.901m/s$. We didn’t find a correlation between KS and $\boldsymbol{p}$; the maximum KS is 0.11 for $G_{z}=5.826K/\mu m,R_{z}=1.165m/s$. Fig. 9d shows MR as a function of the height $z_{l}$. At the height $z_{l}=50\mu m$, the mean and standard deviation are 15.3%$\pm$3.8for 100 $\boldsymbol{p}$. The total number of grain elimination events across the 100 cases is 6,597; the number of true positives predicted by GrainGNN is 6,088.
@@ 3.3 Out-of-distribution generalization
Domain width $L_{x},L_{y}$: A larger domain width with the same grain size distribution is equivalent to a larger number of grains. As discussed in Section 2.3, the dimensional features in Eq. 2.5 and Eq. 2.4, for example $x_{j},y_{j},s_{g},v_{g}$ are normalized by the training domain size $L^{0}_{x},L^{0}_{y},L^{0}_{z}$. Although we change the domain size, we still normalize with $L^{0}$. As shown in Fig. 10a, we have a testing case with domain width $L_{x}=L_{y}=120\mu m$. After normalization, the junction coordinates are in the range of [0, 3]. As the transformer encodes relative distances between a vertex and its neighbors, adding offsets to the coordinates won’t change the hidden states output by the encoder. Thus, we expect no difficulties of using a large domain. We simply evolve GrainGNN on the entire graph and reconstruct the microstructure from the evolved graph.
@@
Fig. 10a shows an example of GrainGNN prediction for a case with domain size $(120\mu m,120\mu m,50\mu m)$. $G_{z}=10K/\mu m,R_{z}=2m/s$. The initial number of grains is 1043. At the end of the simulation, the number of eliminated grains is 704, of which 644 grains are predicted by GrainGNN. The MR at the top layer is 18.2% and the KS statistic for the grain size distribution is 0.021. We further test 20 randomly sampled pairs of $(G,R)$ with domain size $(80\mu m,80\mu m,50\mu m)$, and repeat for domain size $(120\mu m,120\mu m,50\mu m)$. Their MR statistics are shown in Fig. 9f and Fig. 9e. At $z_{l}=50\mu m$, their average MR and standard deviation are 14.4%$\pm$4.0and 13.9%$\pm$3.6%, respectively. Compared to 15.3%$\pm$3.8for in-distribution generalization (Fig. 9d), we don’t observe a noticeable increase in pointwise error when increasing the domain width. The accuracy of accumulative grain events is 92.1% (4,271/4,638) for $L_{x}=2L^{0}_{x}$ and 92.6% (8,956/9,672) for $L_{x}=3L^{0}_{x}$, which is close to 92.3% for $L_{x}=L^{0}_{x}$. In Fig. 10c, we showcase a GrainGNN inference with a large domain width $L_{x}=L_{y}=10L^{0}_{x}$. The total number of grains is 11,600. The number of vertex features of the graph is 313K. The final grain microstructure only requires 15 iterations of the graph, which takes about 220 seconds on one CPU. If we run a phase field simulation with the same configuration, the total number of grid points is is $5000\times 5000\times 100=2.5$ billion. At each grid point if we store five phase field variables and their grain indices, the storage required is 25 billion numbers. The number of times steps for the phase field solve is 50K.
@@
Domain height $L_{z}$: In practice he domain $L_{z}$ is determined by the meltpool depth. Here we test the ability of GrainGNN to generalize for $L_{z}>L^{0}_{z}$ by simply taking more iterations. Recall that we used $L^{0}_{z}=50\mu m$ because this is a typical meltpool depth for the rapid solidification conditions we are modeling. We assume for $z_{l}>L^{0}_{z}$, the graph evolution follows the same pattern as the graph changes from height $L^{0}_{z}-\Delta z_{l}$ to $L^{0}_{z}$. In the network inference, we simply set the constraint $\text{max}(z_{l}/L^{0}_{z})=1$. As shown in Fig. 9f, we extend the predictions of the 20 runs with width $L_{x}=2L^{0}_{x}$ to the height $z_{l}=98\mu m$. From $z_{l}=50\mu m$ to $z_{l}=98\mu m$, MR increases from 14.4%$\pm$4.0to 26.6%$\pm$11.6and the grain elimination accuracy drops from 92.1% to 90.6% (5,729/6,323). We can see from 50 to 90 $\mu m$ MR increases almost linearly with the distance the SLI traveled. From 90 to 98 $\mu m$, the jump of the MR is due to the drastic increase of error for a case with $G_{z}=1.774$ and $R_{z}=1.471$, whose MR reaches 71.5% at $z=98\mu m$. Fig. 10b shows one case with $G_{z}=4.827K/\mu m,R_{z}=0.690m/s$. The KS of grain size distribution is 0.024 for $z_{l}=50\mu m$ and 0.028 for $z_{l}=98\mu m$.
@@
Initial grain size and orientation: We set $L=(80\mu m,80\mu m,50\mu m)\neq L^{0}$ and also vary the initial grain size and orientation distributions. Fig. 11 shows the initial ($z=0$) and final ($z=L_{z}$) grain distributions under $G_{z}=4K/\mu m$ and $R_{z}=0.8m/s$; we also test another four $\boldsymbol{p}$ with the values in Fig. 8. For each $\boldsymbol{p}$, we test ten different size distributions with mean ranging from $d_{0}=2\mu m$ (Fig. 11a) to $d_{0}=7.4\mu m$ (Fig. 11b). Their MR and KS are shown in Fig. 9g where each point is averaged across five different $\boldsymbol{p}$. For $d_{0}$ in the 4-6 $\mu m$ range, MR and KS are close to in-distribution errors. MR is higher for runs with smaller grains and is almost doubled for $d_{0}=2\mu m$. One reason is smaller grains have a higher surface area-to-volume ratio thus higher percentage of pixels will update their grain indices under the same physical parameters. Another reason is smaller $d_{0}$ has higher grain and edge event densities thus introducing larger errors. For larger grains $d_{0}>6\mu m$, we observe higher variance for different $\boldsymbol{p}$. For $G_{z}=1K/\mu m$ and $R_{z}=1.6m/s$, MR increases when $d_{0}>6\mu m$ while MR decreases for the other four $\boldsymbol{p}$. The overall MR for all $d_{0}$ is 13.9% and grain elimination accuracy is 92.3% (14,590/15,801).
We . In the modified orientation distribution, we select dominant grain orientation for $\boldsymbol{\xi}$ in angle $\theta_{0}$ with respect to the $z$-axis. We vary $\theta_{0}$ from 0 to $\pi/4$ with ten values sampled for each $\boldsymbol{p}$ and we sample the five $\boldsymbol{p}$ as discussed for grain size distribution. Fig. 11c and Fig. 11d show the initial orientation distribution with $\theta_{0}=0$ and $\theta_{0}=\pi/4$ respectively. As shown in Fig. 9h, MR is slightly higher for larger $\theta_{0}$, which is associated with more grain eliminations we observed for misaligned grains. KS is almost the same for different $\theta_{0}$. The overall MR and grain elimination accuracy are 12.0% and 91.0% (10,529/11,569), respectively.
@@ 4.1 Accuracy
We compare GrainGNN’s accuracy with three other models, (i) graph convolution embedded LSTM (GC-LSTM) [48], (ii) a graph transformer operator (TransformerConv) [56], and (iii) a graph convolutional operator (GraphConv) [43]. We use the same training data and 2.4 million weights for all models. The metrics we used are RRMSE for $|\Delta\boldsymbol{x}|=\sqrt{\Delta x^{2}+\Delta y^{2}}$ and $\Delta s$, AUC for both events, and the average MR for the 100 testing simulations. Compared to GC-LSTM, GrainGNN adds the attention coefficient $\beta_{i,k}$ in Section 2.5. As shown in Table 1, GrainGNN outperforms GC-LSTM for all the metrics. The self-attention mechanism of GrainGNN is important to learn the spatial correlations between different nodes. If we drop LSTM and only keep the TransformerConv as an encoder, regression accuracy RRMSE-$|\Delta\boldsymbol{x}|$ and RRMSE-$\Delta s$ decrease significantly. GraphConv has a higher AUC-G but lower RRMSE-$|\Delta\boldsymbol{x}|$, RRMSE-$\Delta s$, and AUC-E compared to TransformerConv. We also compute the average MR of 100 testing simulations for different models and GrainGNN produces the most accurate microstructure images.
@@
The ordering of grain and edge events when updating the graph is another factor that affects the accuracy of the prediction. We currently choose the sorting method based on the classifier probabilities; we have not tested alternatives. One constraint of using Algorithm 1 is the input graph can only have triple junctions and the output graph is guaranteed to have only triple junctions. Algorithm 1 has two noteworthy limitations, although they do not seem to limit the overall GrainGNN accuracy. One is the approximation of new coordinates of $j_{1},j_{2}$ in an edge event. The other is GrainGNN doesn’t guarantee planar graph outputs, which means non-physical edge intersections are possible. The intersections could cause grains to have overlapping areas and create holes in the reconstructed images.

### 31 Jan 2024 v2

@@ 3.2 In-distribution generalization
By in-distribution generalization we refer to GrainGNN inference for $L=L^{0}$, unseen (during training) $\boldsymbol{p}$ values, and unseen $\boldsymbol{\xi}$ values sampled by the same procedure we used for training. We show an example of test case in Fig. 8 with $\boldsymbol{p}=(1.904\ K/\mu m,0.558\ m/s)$. Using nearest neighbor interpolation in $\boldsymbol{p}$ space we determine that GrainGNN should be used with $\Delta z=2.4\mu m$, which determines the number of GrainGNN steps. Fig. 8a shows the grain microstructure at four different heights. At each height $z_{l}$, we show the $G_{l}$ and $I_{l}$ with $500\times 500$ spatial resolution. We compare the reconstructed image with the phase field data at the same height. The misclassified grid points are marked in dark red. The error images indicate that the GrainGNN predictions are accurate representations of the microstructure obtained by our phase field simulations. The last two columns compare the evolution of the 3D microstructure for GrainGNN and phase field predictions. We select two grains to illustrate the grain shape evolution. Fig. 8b-e are different measures of GrainGNN accuracy. Fig. 8b shows the cumulative grain elimination events when the SLI reaches different heights. The blue, red, and dashed red lines depict grain eliminations that happened in the phase field simulation only, GrainGNN inference only, and in both, respectively. We can see that the number of eliminations predicted by GrainGNN is overall accurate but lagged after $z_{l}=24\ \mu m$ compared to the phase field result. At $z_{l}=50\ \mu m$, three grains that should be eliminated still exist on the graph, and only one grain is falsely eliminated by GrainGNN. The precision and recall for grain events are 72/73 and 72/75 respectively. Fig. 8c plots MR for reconstructed images at different heights. MR is 2.4% initially at $z_{0}=2\ \mu m$ because the curvature of the grains is neglected in reconstructed images. MR is 11.3% at $z_{l}=50\ \mu m$. Fig. 8d shows the evolution of the volume-averaged misorientation as the height of the SLI increases; the curve is well captured by GrainGNN. The initial average misorientation angle is $23.8^{\circ}$ and it decreases to $13.8^{\circ}$, which indicates the grains are more aligned with the temperature gradients after the epitaxial growth. Fig. 8e shows the grain size distribution when the SLI reaches $50\mu m$. The final average grain size is $9.9\mu m$. The KS statistic between the phase field and GrainGNN distributions is 0.034, which means the two distributions are nearly identical.
@@
Fig. 10: Error statistics for in-distribution and out-of-distribution generalization. (a) Black lines are $\boldsymbol{p}$ grid used for training and validation. Black dots are 100 $\boldsymbol{p}$ values for testing. (b) Misclassification rate (MR) averaged across sampled heights $z_{l}$ for each testing $\boldsymbol{p}$. (c) KS statistics (KS) of grain size distribution at the end of simulations for testing $\boldsymbol{p}$. (d) MR mean and standard deviation for 100 testing $\boldsymbol{p}$ with domain size $(40\mu m,40\mu m,50\mu m)$. (e) MR for domain size $(120\mu m,120\mu m,50\mu m)$. (f) MR for domain size $(80\mu m,80\mu m,100\mu m)$. (g) MR and KS for different initial mean grain sizes $d_{0}$. (h) MR and KS for different initial grain orientation distributions. $\theta_{0}$ is the most frequent misorientation angle between the grain orientation and the $z$-axis.
@@
We further test GrainGNN’s MR accuracy on 100 randomly selected $\boldsymbol{p}$ with a single random $\boldsymbol{\xi}$ for each value of $\boldsymbol{p}$ (see Fig. 10a ). Fig. 10b show the MR averaged over $z_{l}$ for each testing case. MR increases with increasing $R_{z}$ and decreasing $G_{z}$; the highest MR is 17% for $G_{z}=2.450K/\mu m,R_{z}=1.753m/s$ and $G_{z}=3.764K/\mu m,R_{z}=1.901m/s$. We didn’t find a correlation between KS and $\boldsymbol{p}$; the maximum KS is 0.11 for $G_{z}=5.826K/\mu m,R_{z}=1.165m/s$. Fig. 10d shows MR as a function of the height $z_{l}$. At the height $z_{l}=50\mu m$, the mean and standard deviation are 15.3%$\pm$3.8for 100 $\boldsymbol{p}$. The total number of grain elimination events across the 100 cases is 6,597; the number of true positives predicted by GrainGNN is 6,088; the number of false positives is 124.
@@ 3.3 Out-of-distribution generalization
Domain width $L_{x},L_{y}$: A larger domain width with the same grain size distribution is equivalent to a larger number of grains. As discussed in Section 2.3, the dimensional features in Eq. 2.5 and Eq. 2.4, for example $x_{j},y_{j},s_{g},v_{g}$ are normalized by the training domain size $L^{0}_{x},L^{0}_{y},L^{0}_{z}$. Although we change the domain size, we still normalize with $L^{0}$. As shown in Fig. 11a, we have a testing case with domain width $L_{x}=L_{y}=120\mu m$. After normalization, the junction coordinates are in the range of [0, 3]. As the transformer encodes relative distances between a vertex and its neighbors, adding offsets to the coordinates won’t change the hidden states output by the encoder. Thus, we expect no difficulties in using a large domain. We simply evolve GrainGNN on the entire graph and reconstruct the microstructure from the evolved graph.
@@
Fig. 11a shows an example of GrainGNN prediction for a case with domain size $(120\mu m,120\mu m,50\mu m)$. $G_{z}=10K/\mu m,R_{z}=2m/s$. The initial number of grains is 1043. At the end of the simulation, the number of eliminated grains is 704, of which 644 grains are predicted by GrainGNN. The MR at the top layer is 18.2% and the KS statistic for the grain size distribution is 0.021. We further test 20 randomly sampled pairs of $(G,R)$ with domain size $(80\mu m,80\mu m,50\mu m)$, and repeat for domain size $(120\mu m,120\mu m,50\mu m)$. Their MR statistics are shown in Fig. 10f and Fig. 10e. At $z_{l}=50\mu m$, their average MR and standard deviation are 14.4%$\pm$4.0and 13.9%$\pm$3.6%, respectively. Compared to 15.3%$\pm$3.8for in-distribution generalization (Fig. 10d), we don’t observe a noticeable increase in pointwise error when increasing the domain width. The accuracy of accumulative grain events is 92.1% (4,271/4,638) for $L_{x}=2L^{0}_{x}$ and 92.6% (8,956/9,672) for $L_{x}=3L^{0}_{x}$, which is close to 92.3% for $L_{x}=L^{0}_{x}$. In Fig. 11c, we showcase a GrainGNN inference with a large domain width $L_{x}=L_{y}=10L^{0}_{x}$. The total number of grains is 11,600. The number of vertex features of the graph is 313K. The final grain microstructure only requires 15 iterations of the graph, which takes about 220 seconds on one CPU. If we run a phase field simulation with the same configuration, the total number of grid points is $5000\times 5000\times 100=2.5$ billion. At each grid point if we store five phase field variables and their grain indices, the storage required is 25 billion numbers. The number of time steps for the phase field solver is 50K.
@@
Domain height $L_{z}$: In practice the domain height $L_{z}$ is determined by the meltpool depth. Here we test the ability of GrainGNN to generalize for $L_{z}>L^{0}_{z}$ by simply taking more iterations. Recall that we used $L^{0}_{z}=50\mu m$. We assume for $z_{l}>L^{0}_{z}$, the graph evolution follows the same pattern as the previous graph step. In the network inference when $z_{l}>L^{0}_{z}$, we set $z_{l}$ as $L^{0}_{z}$ in Eq. 2.4 and Eq. 2.5. As shown in Fig. 10f, we extend the predictions of the 20 runs with width $L_{x}=2L^{0}_{x}$ to the height $z_{l}=98\mu m$. From $z_{l}=50\mu m$ to $z_{l}=98\mu m$, MR increases from 14.4%$\pm$4.0to 26.6%$\pm$11.6and the grain elimination accuracy drops from 92.1% to 90.6% (5,729/6,323). We can see from 50 to 90 $\mu m$ MR increases almost linearly with the distance the SLI traveled. From 90 to 98 $\mu m$, the jump of the MR is due to the drastic increase of error for a case with $G_{z}=1.774$ and $R_{z}=1.471$, whose MR reaches 71.5% at $z=98\mu m$. Fig. 11b shows one case with $G_{z}=4.827K/\mu m,R_{z}=0.690m/s$. The KS of grain size distribution is 0.024 for $z_{l}=50\mu m$ and 0.028 for $z_{l}=98\mu m$.
@@
Initial grain size and orientation: We set $L=(80\mu m,80\mu m,50\mu m)\neq L^{0}$ and also vary $\boldsymbol{\xi}$. As mentioned in Section 3.1, the initial junction coordinates are randomly selected from [0, 1]. We change the average grain size by varying the number of sampled junctions. Fig. 12 shows the initial ($z=0$) and final ($z=L_{z}$) grain distributions under $G_{z}=4K/\mu m$ and $R_{z}=0.8m/s$; we also test another four $\boldsymbol{p}$ with the values in Fig. 9. For each $\boldsymbol{p}$, we test ten different size distributions with mean ranging from $d_{0}=2\mu m$ (Fig. 12a) to $d_{0}=7.4\mu m$ (Fig. 12b). Their MR and KS are shown in Fig. 10g where each point is averaged across five different $\boldsymbol{p}$. For $d_{0}$ in the 4-6 $\mu m$ range, MR and KS are close to in-distribution errors. MR is higher for runs with smaller grains and is almost doubled for $d_{0}=2\mu m$. One reason is smaller grains have a higher surface area-to-volume ratio thus higher percentage of pixels will update their grain indices under the same physical parameters. Another reason is smaller $d_{0}$ has higher grain and edge event densities thus introducing larger errors. For larger grains $d_{0}>6\mu m$, we observe higher variance for different $\boldsymbol{p}$. For $G_{z}=1K/\mu m$ and $R_{z}=1.6m/s$, MR increases when $d_{0}>6\mu m$ while MR decreases for the other four $\boldsymbol{p}$. The overall MR for all $d_{0}$ is 13.9% and grain elimination accuracy is 92.3% (14,590/15,801).
In the modified orientation distribution, we select dominant grain misorientation angle $\theta_{0}$ with respect to the $z$-axis. We vary $\theta_{0}$ from 0 to $\pi/4$ with ten values sampled for each $\boldsymbol{p}$ and we sample the five $\boldsymbol{p}$ as discussed for grain size distribution. Fig. 12c and Fig. 12d show the initial orientation distribution with $\theta_{0}=0$ and $\theta_{0}=\pi/4$ respectively. As shown in Fig. 10h, MR is slightly higher for larger $\theta_{0}$, which is associated with more grain eliminations we observed for misaligned grains. KS is almost the same for different $\theta_{0}$. The overall MR and grain elimination accuracy are 12.0% and 91.0% (10,529/11,569), respectively.
@@ 4.1 Accuracy
We compare GrainGNN’s accuracy with three other models, (i) a graph convolution LSTM (GC-LSTM) [48], (ii) a graph transformer operator (TransformerConv) [57], and (iii) a graph convolutional operator (GraphConv) [43]. We use the same training data and 2.4 million weights for all models. The metrics we used are RRMSE for $|\Delta\boldsymbol{x}|=\sqrt{\Delta x^{2}+\Delta y^{2}}$ and $\Delta s$, AUC for both events, and the average MR for the 100 testing simulations. Compared to GC-LSTM, GrainGNN adds the attention coefficient $\beta_{i,k}$ in Section 2.5. As shown in Table 2, GrainGNN outperforms GC-LSTM for all the metrics. The self-attention mechanism of GrainGNN is important to learn the spatial correlations between different nodes. If we drop LSTM and only keep the TransformerConv as an encoder, regression accuracy RRMSE-$|\Delta\boldsymbol{x}|$ and RRMSE-$\Delta s$ decrease significantly. GraphConv has a higher AUC-G but lower RRMSE-$|\Delta\boldsymbol{x}|$, RRMSE-$\Delta s$, and AUC-E compared to TransformerConv. We also compute the average MR of 100 testing simulations for different models and GrainGNN produces the most accurate microstructure images.
@@
The ordering of grain and edge events when updating the graph is another factor that affects the accuracy of the prediction. We currently choose the sorting method based on the classifier probabilities; we have not tested alternatives. One constraint of using Algorithm 1 is the input graph can only have triple junctions and the output graph is guaranteed to have only triple junctions. Algorithm 1 has two noteworthy limitations, although they do not seem to limit the overall GrainGNN accuracy. One is the approximation of new coordinates of $j_{1},j_{2}$ in an edge event. The other is GrainGNN doesn’t guarantee planar graph outputs, which means non-physical edge intersections are possible. The intersections could cause grains to have overlapping areas or create holes in the reconstructed images. We currently ignore this issue since it won’t break Algorithm 1 or calculations of quantities of interest.